Find x if the average of 20, 20, 19, 13, and x is 16.
Answer by JulietG(1812)

20+20+19+13+x = 16*5 (five numbers)
72+x = 80
Subtract 72 from each side
x = 8
.
Prove it.
20+20+19+13+8=80
80/5 = 16
Success!